# When Text Embedding Meets Large Language Model: A Comprehensive Survey

### v1

@@ V Text Embedding Understanding with LLMs
V-A Methods for Long Context Compression
Long Context Compression proposes that the input for conditioning a language model can be condensed into a smaller, specifically chosen set of words or weights. These methods reduce context length, leading to faster inference speeds and smaller storage requirements. Based on the form of compressed context, this method can be classified into two categories: soft prompt V-A1 and context distillation V-A2. It is worth noting that in natural language processing, handling long contexts has always been a significant issue. Some works [196, 197, 198, 199, 200, 201, 202] enhance the transformer’s ability to model long contexts by storing and utilizing past representations. These works primarily aim to expand the context of Transformers rather than compressing contexts, but their methods might also be adapted for context compression in the future.
V-A1 Soft Prompt
@@
The restoration method uses text reconstruction in autoencoders as the optimization objective.  [81] utilizes an encoder-decoder structure and trains memory tokens in an reconstruction manner. The original context is converted into shorter memory token embeddings, which are then used to reconstruct the context. [183] employs a set of learnable digest embeddings to condense contextual information, producing digest vectors. [184] utilizes an encoder to segment long contexts into fixed-length embeddings, which are then input into the LLM along with prompts. Then, an instruction reconstruction task is proposed to rebuild the instructions and generate answers. [182] retains key information and compresses other contents into shorter soft prompts. PE-RANK [185] compresses the embedding of each passage into a passage representation. It employs a two-stage ranking training method: in the first stage, reconstruction loss is used to enable the LLM to reconstruct the corresponding paragraph from the input embeddings; in the second stage, ranking loss is applied to equip the LLM with ranking capabilities.
V-A2 Context Distillation
@@
V-B Method for Embedding Inversion
@@
Unlike attribute inference attacks and embedding inversion attacks, the target of model inversion attacks is not the input text but the text embedding model itself. [190] proposed Text Revealer as a model inversion attack for text reconstruction against text classification with transformers. Based on an external dataset and the black-box access to the original text embedding model, it constructed an (embedding, text) training dataset, which is used to train a GPT-2 [215] as a text generator. Using this text generator and perturbing the hidden state optimally with the feedback of the target model, the text revealer could reconstruct private texts in the original training data.  [191] proposes a transfer attack method that employs a surrogate model to emulate the victim model, aiming to steal the text encoder from the victim model. And this paper also uses the GPT-base model as an embedding-to-text part to invert embeddings to their original text sequences.
VI Challenge
VI-A Existing challenges
VI-A1 False Negative detection
Under the embedding training framework of contrastive learning, the model aims to learn appropriate embedding representations by pulling positive sample pairs closer while pushing negative sample pairs farther apart. The appearance of false negatives disrupts this learning process. Samples that originally belong to the same category (positive sample relationship) are wrongly regarded as negative samples to participate in the training. This causes the model to receive incorrect ”supervisory signals”, resulting in the learned embeddings failing to accurately reflect the real semantic relationships among samples. An intuitive idea to solve the false negative problem is to correct the wrongly labeled samples through accurate annotations. However, whether it is the traditional manual annotation method or annotation with the help of large-scale language models (LLMs), both face the challenge of high costs. For manual annotation, in the face of massive data, it is almost an impossible task to expect to annotate thousands of samples for each query. Manual annotation is not only time-consuming and labor-intensive but also easily affected by subjective factors of annotators and issues such as the consistency of annotation standards. On the other hand, although LLMs have demonstrated powerful capabilities in many fields such as natural language processing, using them for annotation also comes with relatively high cost expenditures, including the cost of calling APIs, possible errors in generated results, and limitations on the accuracy of annotating domain-specific knowledge. Therefore, relying on large-scale and accurate annotations to completely eliminate false negatives has significant obstacles in reality.
@@ VI-A1 False Negative detection
VI-A2 The Curse of Low-Resource Languages
@@
VI-A3 Native High-Quality Embedding
@@
VI-B New challenges
VI-B1 Complex Instruction following
@@
Additional, although substantial progress has been made in text representation methods based on Instruction Tuning, it is challenging for users to clearly and accurately describe their intent in practical applications. Therefore, automatically generating clear descriptions based on available information, such as the type of corpus, the user’s query, and other background information, represents a direction that requires further exploration.
VI-B2 Privacy Leakage from Text Embedding
@@
In fact, we can see the antagonism in the two tasks elaborated in Section V: The methods in long context compression aim to compress long text losslessly, while in embedding inversion, the works warn that the information embedded in the embedding can be easily restored and try to design some defenses.
VI-B3 High-Dimensional Text Embedding
Scaling the model size but fixing the embedding dimension has been shown to be an effective way to improve the encoder’s performance [118]. However, increasing the dimension of hidden states with layer and attention head number has been a generalized method when scaling LLMs’ size. Therefore, LLMs’ hidden states generally have a dimension over 2,048, which is 2x and 2.67x than that of BERT-Large and BERT-base, separately. When large models are used as encoders, the representation dimensions exacerbate the computational and storage overhead of the representations. Some post-processing approaches [245, 246, 247] have been proposed to reduce the dimensionality of the learned representations, while matryoshka representation learning (MRL) [172] use the special optimization objectives during training that direct the most critical information to the part dimensions of the representation. The former has not yet been applied to large models, while the latter has already landed in commercial services  [^12] https://openai.com/index/new-embedding-models-and-api-updates. It seems that low-dimensional embedding can obtain a no-bad performance on the MTEB Benchmark. However, there are many deeper issues to be explored. For example, (1) what is the theoretical lower bound on the dimensionality of lossless compression; (2) how to design more efficient compression algorithms to obtain embeddings with higher information density; and (3) how to derive sparse representations that are more friendly to long documents. However, dimension, as one of the most important measures of embedding quality, still has many deeper issues to be explored. For example, (1) what is the theoretical lower bound of dimensionality for information lossless compression of data; (2) how to realize the efficient conversion of low information density high-dimensional embeddings to high-information density low-dimensional embeddings; and (3) how to convert text embedding to sparse representations that are more friendly to long documents.
VI-B4 Training & Inference Overhead for LLMs
@@
VII Future Trends
VII-A Methods for Cross-Lingual & Cross-Modal Domain
@@
VII-A1 Cross-Lingual Representation Learning with LLMs
@@
In cross-lingual scenarios, the downstream tasks, such as STS [43, 46] and retrieval [256, 257], have sufficient high-quality datasets for evaluation. With the help of multilingual PLMs models (e.g., mBERT [9] and XLM-R [9]), mDPR [258], mE5 [259], and M3-Embedding [58] have achieved better results on various types of downstream tasks.
@@
Some LLMs are pre-trained on large multi-lingual corpora, which leads us to expect higher-quality cross-language textual representations using LLMs. Some exploratory work has attempted to introduce LLM into cross-language scenarios and has shown remarkable potential. E5-Mistral [102] demonstrates the great potential of LLM for synthesizing data and being a backbone, while UDEVER [132] shows the powerful generalization capabilities of multilingual LLM, even when fine-tuned using only English text. In addition, both OpenAI and Cohere offer multilingual embedding services; Cohere blogged about the second phase of its model training using additional supervised signals generated by LLM  [^13] https://cohere.com/blog/introducing-embed-v3; and OpenAI’s currently disclosed tech line [131] and up to 3,072 embedding dimensions give us a reasonable suspicion that the LLM backs the service as an encoder.
VII-A2 Cross-Modal Representation Learning with LLMs
@@
In the cross-modal domain, joint language-vision representation learning [260, 261, 262] has received the most attention. Its applications have gradually expanded from image-text retrieval [263, 264] to multiple tasks such as composed image retrieval [265, 266], multi-modal document retrieval [267] and multi-modal knowledge retrieval [268, 269]. VISTA [270] generates high-quality instructions and captions using LLMs and achieves substantial improvements on various cross-modal retrieval tasks, while E5-V [271] demonstrates the potential of MLLMs as a multi-modal unified encoder, showing strong performance on uni-modal (even visual modal) and multi-modal by training the model on text pairs alone.
VII-B Task-Specific Representation
@@
For a long time, text representation learning has been expected to obtain a universal representation without making the downstream task visible. However, this setting is often difficult in practice, such as (1) In the unsupervised setting, contrastive learning brings the two views of data augmentation for the same instance closer together, i.e., learning an encoder that is insensitive to a particular data augmentation transform. However, this may result in valuable information for downstream tasks not being included in the representation [272]; (2) In the supervised setting, undifferentiated joint training of different tasks will hurt performance [273, 104]; (3) In evaluation, the two texts can be evaluated from various perspectives and show different semantic similarities [274, 275];
@@
Therefore, we can convert the hope from universal text representations to universal text encoders $F$, which can be expressed as
@@
The universal encoders accept the text input $x$ and the downstream task $t$ and represent $x$ specific to task $t$, where $t$ can be an identifier or a paragraph description of the task requirements. According to where $t$ acts, we can categorize the methods into pre-processing (also called instruction-following representation) and post-processing (also called equivariant representation learning).
VII-B1 Instruction-Following Representation Learning
In instruction-following representation learning, $F$ is divided into two parts: $F=f\circ p_{t}$, where $p_{t}$ is a prompt template related to task $t$. In the beginning, prompts are used to align the fine-tuning targets with the pre-trained targets of the PLMs [276]. Then, similar instruction tuning [277] is proposed to adjust LLMs to be more adaptable to the format of task-oriented conversations. Instruction-following representation learning follows the idea of instruction tuning but models all tasks into a contrastive paradigm. TART [278] firstly validate the feasibility of this idea to information retrieval, while Instructor [104] fine-tuning GTR on 330 NLP tasks and find that including instructions would alleviate conflicts between tasks compared to traditional multi-task learning. In the zero-shot evaluation, the model fine-tuned with instructions shows good generalization to new instructions and datasets. The findings related to this technique are still iterating: [130] add instruction only on evaluation and demonstrate its effectiveness; [133] demonstrate that question-and-answer pairs can significantly improve a model’s ability to comply with instructions under a particular training method.
VII-B2 Equivariant Representation Learning
In equivariant representation learning, $F$ is divided into two parts: $F=u_{t}\circ f$, where $u_{t}$ is a mapping related to task $t$ in the representation space. Before introducing specific methods, we define “equivariant mapping” in mathematics:
@@ VII-B2 Equivariant Representation Learning
(Equivariance Mapping) Let $f:X\rightarrow\mathbb{R}^{d}$ is a mapping from data to representations. We call $f$ is equivariant to the algebraic group $T=(\mathbb{T},\circ)$ if there exist a transformation $g:\mathbb{T}\times\mathbb{X}\rightarrow\mathbb{X}$ and a transformation $G:\mathbb{T}\times\mathbb{Z}\rightarrow\mathbb{Z}$ satisfying $G_{t}(f(x))=f(g_{t}(x))$ for any $t\in\mathbb{T}$ and $x\in\mathbb{X}$.
$T$ is no longer restricted to algebraic groups in equivariant representation learning. In practice, $t$ is viewed as a data augmentation strategy in the unsupervised setting or a mapping from data to data/label for a specific task in the supervised setting. Specifically, when $G_{t}$ is the identity mapping, the condition in Definition 1 degenerates to $f(x)=f(g_{t}(x))$, i.e., the alignment in contrastive learning [146]. Instance-level equivariant representation learning has been fully explored in self-supervised visual representation learning [279, 280, 281, 282]. In text representation learning, equivariant representation learning based on PLMs is applied in both task-level [273] and instance-level [283, 284, 285].
VII-C Interpretable Text Embedding
Current text representations are far superior at the performance level to those obtained using feature engineering; however, they severely lack interpretability. This leads to an inability to localize why they fail in some cases (e.g., bad cases in information retrieval) to the point where we can’t target them for improvement. Some recent approaches attempt to enhance the interpretability of representations using LLMs.
@@ VII-C Interpretable Text Embedding
[193] tries to use LLM interpretation tables to represent user portraits and movie information by arbitrary representations, even if these representations are constructed by human interpolation; [195] learn the representation for the novel concept to explain them with LLMs;
@@
[114] add The sentence [sent] means in a word:‘‘ before the input text and using last pooling to obtain a textual representation. The method allows the textual representation to decode the token that summarizes the semantics of this text. Similarly, [133] trains the LLM using Q&A pairs containing short answers and generates text representations using the last pooling strategy. [154] find that the contrastive text representations produced by LLMs are highly correlated with the key tokens in the input text, independent of the model architecture, training strategy, and embedding method. [286] directly ask LLMs predefined questions and generates text representations containing only 0-1 from the answers.
VIII Conclusion

### v2

@@ 5 Text Embedding Understanding with LLMs
5.1 Methods for Long Context Compression
Long Context Compression proposes that the input for conditioning a language model can be condensed into a smaller, specifically chosen set of words or weights. These methods reduce context length, leading to faster inference speeds and smaller storage requirements. Based on the form of compressed context, this method can be classified into two categories: soft prompt 5.1.1 and context distillation 5.1.2. It is worth noting that in natural language processing, handling long contexts has always been a significant issue. Some works [247, 248, 249, 250, 251, 252, 253] enhance the transformer’s ability to model long contexts by storing and utilizing past representations.  [254] proposed compressing the raw activations of long contexts. These works primarily aim to expand the context of Transformers rather than compressing contexts, but their methods might also be adapted for context compression in the future.
5.1.1 Soft Prompt
@@
The restoration method uses text reconstruction in autoencoders as the optimization objective.  [132] utilizes an encoder-decoder structure and trains memory tokens in an reconstruction manner. The original context is converted into shorter memory token embeddings, which are then used to reconstruct the context. [234] employs a set of learnable digest embeddings to condense contextual information, producing digest vectors. [235] utilizes an encoder to segment long contexts into fixed-length embeddings, which are then input into the LLM along with prompts. Then, an instruction reconstruction task is proposed to rebuild the instructions and generate answers. [233] retains key information and compresses other contents into shorter soft prompts. PE-RANK [236] compresses the embedding of each passage into a passage representation. It employs a two-stage ranking training method: in the first stage, reconstruction loss is used to enable the LLM to reconstruct the corresponding paragraph from the input embeddings; in the second stage, ranking loss is applied to equip the LLM with ranking capabilities.  [257] proposed fine-grained autoencoding and segment-wise token importance estimation to enhance gist-based context compression.
5.1.2 Context Distillation
@@
5.2 Method for Embedding Inversion
@@
Unlike attribute inference attacks and embedding inversion attacks, the target of model inversion attacks is not the input text but the text embedding model itself. [241] proposed Text Revealer as a model inversion attack for text reconstruction against text classification with transformers. Based on an external dataset and the black-box access to the original text embedding model, it constructed an (embedding, text) training dataset, which is used to train a GPT-2 [267] as a text generator. Using this text generator and perturbing the hidden state optimally with the feedback of the target model, the text revealer could reconstruct private texts in the original training data.  [242] proposes a transfer attack method that employs a surrogate model to emulate the victim model, aiming to steal the text encoder from the victim model. This paper also uses the GPT-base model as an embedding-to-text part to invert embeddings to their original text sequences.
6 Challenge
6.1 Existing challenges
6.1.1 False Negative detection
Under the embedding training framework of contrastive learning, the model aims to learn appropriate embeddings by pulling positive sample pairs closer while pushing negative sample pairs farther apart. The appearance of false negatives disrupts this learning process. Samples that originally belong to the same category (positive sample relationship) are wrongly regarded as negative samples to participate in the training. This causes the model to receive incorrect "supervisory signals", resulting in the learned embeddings failing to accurately reflect the real semantic relationships among samples. An intuitive idea to solve the false negative problem is to correct the wrongly labeled samples through accurate annotations. However, whether it is the traditional manual annotation method or annotation with the help of large-scale language models (LLMs), both face the challenge of high costs. For manual annotation, in the face of massive data, it is almost an impossible task to expect to annotate thousands of samples for each query. Manual annotation is not only time-consuming and labor-intensive but also easily affected by subjective factors of annotators and issues such as the consistency of annotation standards. On the other hand, although LLMs have demonstrated powerful capabilities in many fields such as natural language processing, using them for annotation also comes with relatively high cost expenditures, including the cost of calling APIs, possible errors in generated results, and limitations on the accuracy of annotating domain-specific knowledge. Therefore, relying on large-scale and accurate annotations to completely eliminate false negatives has significant obstacles in reality.
@@ 6.1.1 False Negative detection
6.1.2 The Curse of Low-Resource Languages
@@
6.1.3 Native High-Quality Embedding
@@
6.2 New challenges
6.2.1 Complex Instruction following
@@
Additionally, although substantial progress has been made in text embedding methods based on Instruction Tuning, it is challenging for users to clearly and accurately describe their intent in practical applications. Therefore, automatically generating clear descriptions based on available information, such as the type of corpus, the user’s query, and other background information, represents a direction that requires further exploration.
6.2.2 Privacy Leakage from Text Embedding
@@
In fact, we can see the antagonism in the two tasks elaborated in Section 5: The methods in long context compression aim to compress long text losslessly, while in embedding inversion, the works warn that the information embedded in the embedding can be easily restored and try to design some defenses.
6.2.3 High-Dimensional Text Embedding
Scaling the model size but fixing the embedding dimension has been shown to be an effective way to improve the encoder’s performance [175]. However, increasing the dimension of hidden states with layer and attention head number has been a generalized method when scaling LLMs’ size. Therefore, LLMs’ hidden states generally have a dimension over 2,048, which is 2x and 2.67x than that of BERT-Large and BERT-base, separately. When LLMs are used as embedders, the increased embedding dimensions exacerbate the computational and storage overhead.
To achieve a flexible trade-off between performance and efficiency, Matryoshka representation learning (MRL) [223] uses the special optimization objectives during training that direct the most critical information to the part dimensions of the embedding. This technique has already landed in commercial services  [^18] https://openai.com/index/new-embedding-models-and-api-updates. In addition, some post-processing approaches [293, 294, 295] have been proposed to reduce the dimensionality of the learned embeddings. A recent work, CSR [296], introduces automatic coding and task-aware contrastive learning to learn sparse representations. The authors show the sparse representation can maintain almost the same performance as the original LLM-based Embedder in 32 dimensions, demonstrating the huge potential for post-sparse techniques.
In addition, there are many deeper issues to be explored. For example, (1) what is the theoretical lower bound on the dimensionality of lossless compression; (2) how to design more efficient compression algorithms to obtain embeddings with higher information density; and (3) how to derive sparse representations that are more friendly to long documents. However, dimension, as one of the most important measures of embedding quality, still has many deeper issues to be explored. For example, (1) what is the theoretical lower bound of dimensionality for information lossless compression of data; (2) how to realize the efficient conversion of low information density high-dimensional embeddings to high-information density low-dimensional embeddings; and (3) how to convert text embedding to sparse representations that are more friendly to long documents.
6.2.4 Training & Inference Overhead for LLMs
@@
7 Future Trends
7.1 Methods for Cross-Lingual & Cross-Modal Domain
@@
7.1.1 Cross-Lingual Text Embedding
@@
In cross-lingual scenarios, the downstream tasks, such as STS [110, 117] and retrieval [304, 57], have sufficient high-quality datasets for evaluation. With the help of multilingual PLMs models (e.g., mBERT [68] and XLM-R [68]), mDPR [305], mE5 [306], BGE-M3 [56], mGTE [307] and KaLM-Embedding [308] have achieved better results on various types of downstream tasks.
@@
Some LLMs are pre-trained on large multi-lingual corpora, which leads us to expect higher-quality cross-language text embeddings using LLMs. Some exploratory work has attempted to introduce LLM into cross-language scenarios and has shown remarkable potential. E5-Mistral [153] demonstrates the great potential of LLM for synthesizing data and being a backbone, while UDEVER [190] shows the powerful generalization capabilities of multilingual LLM, even when fine-tuned using only English text.
In addition, both OpenAI and Cohere offer multilingual embedding services; Cohere blogged about the second phase of its model training using additional supervised signals generated by LLM  [^19] https://cohere.com/blog/introducing-embed-v3; and OpenAI’s currently disclosed tech line [189] and up to 3,072 embedding dimensions give us a reasonable suspicion that the LLM backs the service as an encoder.
7.1.2 Cross-Modal Text Embedding
@@
In the cross-modal domain, joint language-vision embedding learning [309, 310, 311] has received the most attention. Its applications have gradually expanded from image-text retrieval [312, 313] to multiple tasks such as composed image retrieval [314, 315], multi-modal document retrieval [316] and multi-modal knowledge retrieval [317, 318]. VISTA [319] generates high-quality instructions and captions using LLMs and achieves substantial improvements on various cross-modal retrieval tasks. For example, E5-V [320] fine-tuning MLLMs with text-modal contrastive learning only, firstly showing the potential of MLLMs as a multi-modal unified encoder; GME [321] synthesized large-scale and high-quality data for fine-tuning MLLM through the pipeline, showing strong performance on uni-modal (even visual modal) and multi-modal by training the model on text pairs alone. LLaVE [322]
7.2 Task-Specific Text Embedding
@@
For a long time, text embedding learning has been expected to obtain a universal embedding without making the downstream task visible. However, this setting is often difficult in practice, such as (1) In the unsupervised setting, contrastive learning brings the two views of data augmentation for the same instance closer together, i.e., learning an encoder that is insensitive to a particular data augmentation transform. However, this may result in valuable information for downstream tasks not being included in the embedding [323]; (2) In the supervised setting, undifferentiated joint training of different tasks will hurt performance [324, 67]; (3) In evaluation, the two texts can be evaluated from various perspectives and show different semantic similarities [325, 326];
@@
Therefore, we can convert the hope from universal text embeddings to universal text encoders $F$, which can be expressed as
@@
The universal encoders accept the text input $x$ and the downstream task $t$ and represent $x$ specific to task $t$, where $t$ can be an identifier or a paragraph description of the task requirements. According to where $t$ acts, we can categorize the methods into pre-processing (also called instruction-following embedding) and post-processing (also called equivariant embedding learning).
7.2.1 Instruction-Following Embedding
In instruction-following embedding learning, $F$ is divided into two parts: $F=f\circ p_{t}$, where $p_{t}$ is a prompt template related to task $t$. In the beginning, prompts are used to align the fine-tuning targets with the pre-trained targets of the PLMs [327]. Then, similar instruction tuning [328] is proposed to adjust LLMs to be more adaptable to the format of task-oriented conversations. Instruction-following embedding learning follows the idea of instruction tuning but models all tasks into a contrastive paradigm. TART [51] firstly validate the feasibility of this idea to information retrieval, while Instructor [67] fine-tuning GTR on 330 NLP tasks and find that including instructions would alleviate conflicts between tasks compared to traditional multi-task learning. In the zero-shot evaluation, the model fine-tuned with instructions shows good generalization to new instructions and datasets. The findings related to this technique are still iterating: [164] add instruction only on evaluation and demonstrate its effectiveness; [191] demonstrate that question-and-answer pairs can significantly improve a model’s ability to comply with instructions under a particular training method.
7.2.2 Equivariant Embedding Learning
In equivariant Embedding learning, $F$ is divided into two parts: $F=u_{t}\circ f$, where $u_{t}$ is a mapping related to task $t$ in the embedding space. Before introducing specific methods, we define “equivariant mapping” in mathematics:
@@ 7.2.2 Equivariant Embedding Learning
(Equivariance Mapping) Let $f:X\rightarrow\mathbb{R}^{d}$ is a mapping from data to embeddings. We call $f$ is equivariant to the algebraic group $T=(\mathbb{T},\circ)$ if there exist a transformation $g:\mathbb{T}\times\mathbb{X}\rightarrow\mathbb{X}$ and a transformation $G:\mathbb{T}\times\mathbb{Z}\rightarrow\mathbb{Z}$ satisfying $G_{t}(f(x))=f(g_{t}(x))$ for any $t\in\mathbb{T}$ and $x\in\mathbb{X}$.
$T$ is no longer restricted to algebraic groups in equivariant embedding learning. In practice, $t$ is viewed as a data augmentation strategy in the unsupervised setting or a mapping from data to data/label for a specific task in the supervised setting. Specifically, when $G_{t}$ is the identity mapping, the condition in Definition 1 degenerates to $f(x)=f(g_{t}(x))$, i.e., the alignment in contrastive learning [199]. Instance-level equivariant embedding learning has been fully explored in self-supervised visual embedding learning [329, 330, 331, 332]. In NLP, equivariant embedding learning based on PLMs is applied in both task-level [324] and instance-level [333, 334, 335].
7.3 Interpretable Text Embedding
Current text embeddings are far superior at the performance level to those obtained using feature engineering; however, they severely lack interpretability. This leads to an inability to localize why they fail in some cases (e.g., bad cases in information retrieval) to the point where we can’t target them for improvement. Some recent approaches attempt to enhance the interpretability of embeddings using LLMs.
@@ 7.3 Interpretable Text Embedding
[244] tries to use LLM interpretation tables to represent user portraits and movie information by arbitrary embeddings, even if these embeddings are constructed by human interpolation; [246] learn the embedding for the novel concept to explain them with LLMs;
@@
[166] add The sentence [sent] means in a word:‘‘ before the input text and using last pooling to obtain a text embedding. The method allows the text embedding to decode the token that summarizes the semantics of this text. Similarly, [191] trains the LLM using Q&A pairs containing short answers and generates text embeddings using the last pooling strategy. [206] find that the contrastive text embeddings produced by LLMs are highly correlated with the key tokens in the input text, independent of the model architecture, training strategy, and embedding method. [336] directly ask LLMs predefined questions and generates text embeddings containing only 0-1 from the answers.
8 Conclusion